OpenStudy (anonymous):

What is the d value in the following arithmetic sequence? -7, -2, 3, 8,....

OpenStudy (anonymous):

-7, -2, 3, 8, 13,1 8, 23, 28 and so on just keep adding 5

OpenStudy (anonymous):

*18 my bad

OpenStudy (anonymous):

what does it mean by d value?

OpenStudy (kirbykirby):

It is the common difference between each term. You can always find it by doing ( one term) - (the previous term) so: -2 - (-7) = 5 or 3 - (-2) = 5 or 18 - 13 = 5 etc.. Which is basically what @Thugchick14 described